OpenStudy (vickyisthesmartone123):

What is the lateral area of the rectangular prism with a base length of 14 m, a base width of 9 m, and a height of 6 m?

OpenStudy (vickyisthesmartone123):

@PizzaLover123

OpenStudy (anonymous):

528 A=2(wl+hl+hw)=2·(9·14+6·14+6·9)=528

OpenStudy (vickyisthesmartone123):

thanks

OpenStudy (anonymous):

Your welcome ;) Medal please!!